Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day08


Crew #108 Log Book for December 10, 2011

Photos of the day





Commander's Check-In Report

Date and time: 10 December 2011

Written by:

Charlotte Poupon, Crew 108, Commander

Crew Physical Status:

Perfect.

Time Departed/Returned from EVA:

EVA # 14 / 09:55 to 11:45


Brief Narrative of Field Mission Results:

Not applicable


EVA Data/Interpretations:

Exploration Condor Chasma


Engineering/Hab Maintenance:

Engineers fixing suits.


Report Transmission Schedule:

All reports by 22:00.


Plans for Tomorrow:

light mesurement to start
next attempt to try the rover


Inventory:

Not applicable


Support Requested:

Not applicable


Miscellaneous:

Commander's Report

Date and time:

10 December 2011


Written by:

Charlotte Poupon, Crew 108, Commander


Today was the mid-rotation and we've planned to make it unforgettable.

We start by a very early waking up to look at the lunar eclipse. Then enjoying a cooking day we did a huge breakfast with scramble eggs, bacon and bread. It as awesome !

Then a team did an EVA (documented separately), while crew members remaining in the Hab tidied the living quarter to go on with a good organization for the coming week.

Lunch was also bright (documented separately).

Testing the rover gave us the opportunity to breath from the pressurized tunnel and to have a nice team building experience.

The mood is brilliant in the sim !

Health and Safety Report

Date and time:

Week 1 - Dec 3 to Dec 10, 2011



Written by:

Michael LeClair,
Crew 108, HSO

Note:

One crew member removed from Hab and Simulation on Wednesday December 7th, 2011.


The physical health of all 5 crew-members is excellent. No incidents or accidents to report.

The mental health of all 5 crew-members is also excellent.

Engineering Report

Date and time:

10 December 2011, 19
:20


Written by:

Mike Lotto, Crew 108, Engineer



Generator/Electricity:

Did you have any electricity outage? (Yes/No): No

Fuel Status:
Diesel (Full/0.75/0.5/0.25/Empty): 0.5
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.75
Notes/Comments:

Oil:
ATV Oil Quantity (# of Quarts in storage): 3
Car Oil Quantity (# of Quarts in storage): 3

Water Status:
Outside Potable Water Tank Level (inches from bottom): 10
Trailer Potable Water Tank Level (inches from bottom): 0.5
Hab Potable Water Tank Level (inches from bottom): 13
Potable Water Meter Reading: Unknown
In to GreenHab Meter Reading: Offline
Notes/Comments:

GreenHab:
Condition of plants in Tank 1: Offline
Condition of plants in Tank 2: Offline
Condition of plants in Tank 3: Offline
Greenhab Notes/Comments: GreenHab is offline this season.

Observatory:
Telescope Used? (Yes/No): No
Observatory Notes/Comments:

Transportation:
Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): Yes
Tire Status: Good
Odometer at end of day: 125455 miles
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: No
Fuel Consumed: 0.3 gallons
Tire Status: Good
Comments:

ATV 2 (Spirit):
Used: No
Oil Checked: No
Fuel Consumed: 0 gallons
Tire Status: Good
Comments:

ATV 3 (Viking I):
Used: Yes
Oil Checked: No
Fuel Consumed: 0.3 gallons
Tire Status: Good
Comments:

Heating and Ventilation:
Thermostat setting upstairs (Farenheit): 65
Actual temperature upstairs (Farenheit): 70.3

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 47

Computers/Networking Infrastructure:
Notes/comments on internet/computers?: None

General Engineering Notes/Comments:

EVA 14 Report

Date and time:

10 December 2011


Written by:

Michael LeClair, Crew 108, Geologist



EVA Duration: 09:52 – 11:48


- Depressurization: 09h52 – 09h57.

- ATV to ridge due west of Hab via Lowell Hwy to northern access road to Cactus road, approaching chasm from the north then parking
- Enroute 10h02 to 10h11.


- Foot traverse down through canyon; return same.

- Halite samples collected. Possible site of interest for Astrobiologist identified.

- Progressive photographs of strata collected throughout.

- ATV east (direct) back to Hab via Cactus road. Enroute 11:36 to 11:42.

- Repressurization: 11:43 – 11:48.


Crew Members:

Michael LeClair (Geologist), Mike Lotto (Engineer)



Site Location:

Candor Chasm



Transit Mode:

ATV 1 & 3



Objectives:

Geological Exploration of Candor Chasma



Results:

Possible eugenic salt-induced fracturing identified in 2 distinct silty/clay strata of the Summerville (?) formation.



Lessons Learned:

Morrison and Summerville Formations better delineated and correlated with existing studies.

Journalist Report

Date and time:

10 December, 19:15



Written by:

Ashley Dale, Crew 108, Engineer



Today started at 05:30, with a lunar eclipse and the closest thing we could get to an English breakfast. Being the mid-way point in the rotation, we decided it would be a good idea to take a break. With only one EVA and a test-run of Buddy Bot (our Mars Rover), and the odd engineering maintenance job, the day was really quite relaxed. We rigged up the sound system and had all sorts playing over the day whilst writing up personal reports, reading books, and generally gelling together. “Symphony of Science – The Case for Mars” was a favorite, naturally.


Buddy Bot Report 1

Date:

10 December 2011


Written by :

Charlotte Poupon, Crew 108 Commander, industrial designer, mother of the BuddyBot

Text

Yesterday, we found out that the BuddyBot cannot work because the battery charger used is not compatible with 110V. We tried to find a solution by using other charger system but it would have been something like fire hazard, so we did not even try.

After a quite worry-plagued night - assuming I was ashamed of this fail - I started to consider using the spare batteries I bought.

As BuddyBot needs 7.2V to work, I connected in series five AAA batteries (5 x 1.5 = 7.5 v) with cello-tape and spare cables, and then this fortune battery with the BuddyBot power cable, which was a totally safe DIY.

AND IT WORKS !

So this afternoon, BuddyBot will make his first wheel turns on the Martian ground.

We'll make a scientific movie from BuddyBot, plus a technical one showing the whole process.

I keep on recommending Mission Support to afford a converter 110 > 220
for future rotations.

Buddy Bot Report 2


Date:

10 December 2011

Written by :

Charlotte Poupon, Crew 108 Commander, industrial designer,  mother of the Buddy Bot

Text

After the huge effort to replace the original uncharged battery, we made
the first try of the robot from the pressurized tunnel, since the rover supposely allows us to operated it from inside the Hab.

Firts, we understood that the bottle meant to stabilized it were too heavy, so we got rid of them but still we had difficulties to operate it on an unstable ground around the Hab.

Buddy Bot is a rear-wheel drive and we found it easier to make it drive backwards which totaly ruin to possibility to operate either navigation camera (with a live remote screen) or the scientific on (a GoPro recording on Memory card).

More changes will be done, to make it lighter and to flip the instruments and other tests will be done by tomorrow.

Chef's Report 1

Date:

10 December 2011

Written by:

Charlotte Poupon, Crew 108, Commander


1a) Today was

- a cooking day

1b) we were celebrating our mid-rotation, so we did have huge meals... this is the report for breakfast + lunch...

2a) List the foods served at today's breakfast, giving full names of each.

- eggs
- bacon
- potatoes
- bread
- jam
- nutella

2b) List the foods served at today's lunch, giving full names of each.

- Quinoa
- green beans
- corn
- butter

3) List any main meal foods not finished by the crew (leftovers stored
for later use or discarded)

None

4) Recipes for anything more complicated than following package
directions exactly, or rehydrating.

N/A

5) Lessons learned

I realized that we talk A LOT about food. After getting in touch though twitter with Romain Charles, the French crew member involved in Mars500 it happens that the marsonauts DID TALK about food all the day as well.

I think it might be very encouraging for your study !


6) Comments/questions for Kim and Jean, the food study investigators

We miss a dish to go on oven (and make some cheesy potatoes baked)...
for brownies we've tried the Pirex one but these ones are not oven suitable (just microwave)... so it... exploded... without injuries... but OK...

Chef's Report 2

Date:

10 December 2011

Written by:

Ashley Dale, Crew 108, Engineer



1a) Today was

- a cooking day


2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

- Tortellini pasta

- Sun-dried tomatoes

- Dehydrated mushrooms (rehydrated and fried)

- Dehydrated broccoli

- Four cheese sauce mix

- Mixed herbs

- Butter

- Milk

- Olive oil


3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

None



4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

N/A


5) Lessons learned

N/A


6) Comments/questions for Kim and Jean, the food study investigators

Could do with replacing the metal grid over the hobs – it is bent towards the middle. Cooking with liquids on a slanted surface isn’t particularly useful.